On January 1 of the current year, the Barton Corporation issued 10% bonds with a face value of $79,000. The bonds are sold for $76,630. The bonds pay interest semiannually on June 30 and December 31 and the maturity date is December 31, five years from now. Barton records straight-line amortization of the bond discount. The bond interest expense for the year ended December 31 is

a.$7,900

b.$8,374

c.$658

d.$2,370
Answer #1

Bond interest payable = 79,000*10%= 7900

Amortization of bond discount

= (79,000-76,630)/5 = 474

Bond interest expense for the year ended Dec 31

= 7900+474

= 8374